find the path of center of mass and center of mass velocity if a chain of length L is moving on a circular path of radius R with velocity v.and L=pi R.
1
Expert's answer
2020-09-10T11:59:18-0400

Explanations & Calculations




  • Consider the figure to get a better understanding.
  • Distance to the com (ox) of an arc of radius r is given by, "\small ox = \large \frac{r \sin \alpha}{\alpha}"
  • Since "\small L= \pi R", the chain makes a semicircular arc whose com lies at a distance from the origin as "\small ox = \large \frac{R \sin \frac{\pi}{2}}{\frac{\pi}{2}} = \frac{2R}{\pi}"


  • Therefore, as the chain rotates, its center of mass also changes it position being fixed at x which describes a circular path of radius ox.


  • Since the chain moves on a circular path at a constant speed, it has an (constant) angular velocity "\small \omega", which is a constant value about the axis for a given circular motion .
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small \omega &= \small \frac{V}{R}= \frac{V_1}{\text{ox}}\\
\small \frac{V}{R}&= \small \frac{V_1}{\frac{2R}{\pi}}\\
\small V_1 &= \small \frac{2V}{\pi}
\end{aligned}"